# Feed Two Birds with One Scone: Exploiting Wild Data for Both Out-of-Distribution Generalization and Detection

### v1

@@ Appendix D Results on Additional OOD Datasets
In this section, we provide the main results on more OOD datasets including Textures (Cimpoi et al., 2014) and LSUN_Resize (Yu et al., 2015) in Table 7. We observe that our proposed approach achieves overall strong performance in OOD generalization and OOD detection on these additional OOD datasets. Particularly, we compare our method with post hoc OOD detection methods such as MSP (Hendrycks & Gimpel, 2017), ODIN (Liang et al., 2018), Energy (Liu et al., 2020), Mahalanobis (Lee et al., 2018), ViM (Wang et al., 2022a), and KNN (Sun et al., 2022). These methods are all based on a model trained with cross-entropy loss, which suffers from limiting OOD generalization performance (75.05%). In contrast, our method achieves an improved OOD generalization performance (e.g., 85.56% when the wild data is a mixture of CIFAR-10, CIFAR-10-C, and Texture).
We also compare our method with common OOD generalization baseline methods including IRM (Arjovsky et al., 2019), Mixup (Zhang et al., 2018), and VREx (Krueger et al., 2021). Our method consistently achieves better results compared to these OOD generalization baselines. Lastly, we compare our method with strong OOD detection methods using $\mathbb{P}_{\text{wild}}$ such as oe (Hendrycks et al., 2018), energy-regularized learning (Liu et al., 2020), and WOODS (Katz-Samuels et al., 2022). Our method demonstrates strong performance on OOD generalization accuracy, which shows the effectiveness of our method for making use of the covariate OOD data.
@@ Appendix E Results on PACS
In this section, we report results on the PACS dataset (Li et al., 2017) and present the comparisons with other baseline methods from the DomainBed. As shown in Table 8, we compare our method with a collection of common OOD generalization baselines, including ERM (Vapnik, 1999), IRM (Arjovsky et al., 2019), GroupDRO (Sagawa et al., 2019), I-Mixup (Zhang et al., 2018), VREx (Krueger et al., 2021), MLDG (Li et al., 2018a), CORAL (Sun & Saenko, 2016), MMD (Li et al., 2018b), DANN (Ganin et al., 2016), CDANN (Li et al., 2018c), MTL (Blanchard et al., 2021), SagNet (Nam et al., 2021), ARM (Zhang et al., 2021), RSC (Huang et al., 2020). Our approach SCONE ($86.4\%$) outperforms all of these OOD generalization baselines on the DomainBed benchmark.
@@
ERM (Vapnik, 1999) 88.1 77.9 97.8 79.1 85.7
IRM (Arjovsky et al., 2019) 85.0 77.6 96.7 78.5 84.4
GroupDRO (Sagawa et al., 2019) 86.4 79.9 98.0 72.1 84.1
I-Mixup (Wang et al., 2020; Xu et al., 2020) 86.5 76.6 97.7 76.5 84.3
VREx (Krueger et al., 2021) 86.0 79.1 96.9 77.7 84.9
MLDG (Li et al., 2018a) 89.1 78.8 97.0 74.4 84.8
CORAL (Sun & Saenko, 2016) 87.7 79.2 97.6 79.4 86.0
MMD (Li et al., 2018b) 84.5 79.7 97.5 78.1 85.0
DANN (Ganin et al., 2016) 85.9 79.9 97.6 75.2 84.6
CDANN (Li et al., 2018c) 84.0 78.5 97.0 71.8 82.8
MTL (Blanchard et al., 2021) 87.5 77.1 96.4 77.3 84.6
SagNet (Nam et al., 2021) 87.4 80.7 97.1 80.0 86.3
ARM (Zhang et al., 2021) 86.8 76.8 97.4 79.3 85.1
RSC (Huang et al., 2020) 85.4 79.7 97.6 78.2 85.2
@@ Appendix F Results on Different Corruption Types
In this section, we provide additional ablation studies of the different covariate shifts. In Table 10, we evaluate our method under 19 different common corruptions such as gaussian noise, defocus blur, glass blur, impulse noise, shot noise, snow, zoom blur, brightness, etc. We follow the default design and parameter setting as in the original paper (Hendrycks & Dietterich, 2018) for generating the corruptions. For our method with margin, $\eta$ is chosen based on the validation strategy in Appendix B. Our method is overall more robust under different covariate shifts than the WOODS baseline.
@@
Covariate shift type Method OOD Acc.normal-$\uparrow$ ID Acc.normal-$\uparrow$ FPRnormal-$\downarrow$ AUROCnormal-$\uparrow$

### v2

@@ Appendix D Results on Additional OOD Datasets
In this section, we provide the main results on more OOD datasets including Textures (cimpoi2014describing) and LSUN_Resize (yu2015lsun) in Table 7. We observe that our proposed approach achieves overall strong performance in OOD generalization and OOD detection on these additional OOD datasets. Particularly, we compare our method with post hoc OOD detection methods such as MSP (hendrycks2016baseline), ODIN (liang2018enhancing), Energy (liu2020energy), Mahalanobis (lee2018simple), ViM (wang2022vim), and KNN (sun2022out). These methods are all based on a model trained with cross-entropy loss, which suffers from limiting OOD generalization performance (75.05%). In contrast, our method achieves an improved OOD generalization performance (e.g., 85.56% when the wild data is a mixture of CIFAR-10, CIFAR-10-C, and Texture).
We also compare our method with common OOD generalization baseline methods including IRM (arjovsky2019invariant), Mixup (zhang2018mixup), and VREx (krueger2021out). Our method consistently achieves better results compared to these OOD generalization baselines. Lastly, we compare our method with strong OOD detection methods using $\mathbb{P}_{\text{wild}}$ such as oe (hendrycks2018deep), energy-regularized learning (liu2020energy), and WOODS (katz2022training). Our method demonstrates strong performance on OOD generalization accuracy, which shows the effectiveness of our method for making use of the covariate OOD data.
@@ Appendix E Results on PACS
In this section, we report results on the PACS dataset (li2017deeper) and present the comparisons with other baseline methods from the DomainBed. As shown in Table 8, we compare our method with a collection of common OOD generalization baselines, including ERM (vapnik1999overview), IRM (arjovsky2019invariant), GroupDRO (sagawa2019distributionally), I-Mixup (zhang2018mixup), VREx (krueger2021out), MLDG (li2018learning), CORAL (sun2016deep), MMD (li2018domain), DANN (ganin2016domain), CDANN (li2018deep), MTL (blanchard2021domain), SagNet (nam2021reducing), ARM (zhang2020adaptive), RSC (huang2020self). Our approach SCONE ($86.4\%$) outperforms all of these OOD generalization baselines on the DomainBed benchmark.
@@
ERM (vapnik1999overview) 88.1 77.9 97.8 79.1 85.7
IRM (arjovsky2019invariant) 85.0 77.6 96.7 78.5 84.4
GroupDRO (sagawa2019distributionally) 86.4 79.9 98.0 72.1 84.1
I-Mixup (wang2020heterogeneous; xu2020adversarial) 86.5 76.6 97.7 76.5 84.3
VREx (krueger2021out) 86.0 79.1 96.9 77.7 84.9
MLDG (li2018learning) 89.1 78.8 97.0 74.4 84.8
CORAL (sun2016deep) 87.7 79.2 97.6 79.4 86.0
MMD (li2018domain) 84.5 79.7 97.5 78.1 85.0
DANN (ganin2016domain) 85.9 79.9 97.6 75.2 84.6
CDANN (li2018deep) 84.0 78.5 97.0 71.8 82.8
MTL (blanchard2021domain) 87.5 77.1 96.4 77.3 84.6
SagNet (nam2021reducing) 87.4 80.7 97.1 80.0 86.3
ARM (zhang2020adaptive) 86.8 76.8 97.4 79.3 85.1
RSC (huang2020self) 85.4 79.7 97.6 78.2 85.2
@@ Appendix F Results on Different Corruption Types
In this section, we provide additional ablation studies of the different covariate shifts. In Table 10, we evaluate our method under 19 different common corruptions such as gaussian noise, defocus blur, glass blur, impulse noise, shot noise, snow, zoom blur, brightness, etc. We follow the default design and parameter setting as in the original paper (hendrycks2018benchmarking) for generating the corruptions. For our method with margin, $\eta$ is chosen based on the validation strategy in Appendix B. Our method is overall more robust under different covariate shifts than the WOODS baseline.
@@
[…]